Banc Corp. Trust is considering either a bankwide overhead rate or department overhead rates to allocate $435,750 of indirect costs. The bankwide rate could be based on either direct labor hours (DLH) or the number of loans processed. The departmental rates would be based on direct labor hours for Consumer Loans and a dual rate based on direct labor hours and the number of loans processed for Commercial Loans. The following information was gathered for the upcoming period:

Department DLH Loans Processed Direct Costs
Consumer 15,000 850 $ 260,000
Commercial 8,000 200 $ 140,000

If Banc Corp. Trust uses a bankwide rate based on the number of loans processed, what would be the total costs for the Commercial Department?
Answer #1

Bankwide rate based on no. of loans processed = Total indirect costs/Total no. of loans processed

Total no. of loans processed = 850+200 = 1,050 loans

Bank wide rate = $435,750/1,050 loans = $415 per loan processed

Indirect cost allocated to Commercial Department = $415 per loan*200 loans processed

= $83,000

Total costs for the Commercial Department = Direct cost+Indirect cost allocated

= $140,000+$83,000 = $223,000

Therefore total costs for the Commercial Department is $223,000.